The following graph describes the motion of Abdul going to meet his friend who stays 50m from his house.a)How much time is required to his friend’s house?b)What is the distance travelled by Abdul during the time interval 0 to 12 min?c)During which time interval is he moving towards his house?d)For how many minutes was he at rest during the entire journey?e)Calculate the speed by which she returned home.
1
Expert's answer
2020-04-20T10:43:42-0400


a)


"t=14\ min"

b)


"d=20+20+20=60\ m"

c)


"\text{from 8 to 10 min and from 12 to 14 min}"

d)


"t=(4-2)+(8-6)+(12-10)=6\ min"

e)


"v=\frac{50}{2}=25\frac{m}{min}"